Tohban Report 2012-12-19

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 12 Dec 2012
End Date: 19 Dec 2012
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Mitsuo Oka (
List all reports


Solar Activity

The Sun remained quiet this week. Currently there are four active regions on the disk. STEREO-B does not show new regions about to rotate on, so activity is likely to stay quiet for the following week.

How many GOES flares occurred?

 Flares above B, C, M, X class were     37     12     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     28     9     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     17     4     0     0

There were RHESSI flares/GOES flares 191 / 49 over the time range 12-Dec-12 11-Dec-19

Memory Management

We began this week with an effort to empty the SSR, which had been filled to a high level last week by broadband noise in detector 8. (The noise was determined to go up to hundreds of keV, effectively hindering any meaningful data-taking.) To this end, D8 events were disabled just prior to the start of this week's tohbanning. This was effective in bringing down the memory fill, which eventually dropped to zero on Sunday.

But two steps forward; one step back: multiple ground stations reported problems this week! The Wallops station often recorded bad data during the time of closest approach of the pass, resulting in a frequent necessity to replay data. They are supposed to have this problem fixed soon. On top of that, a BGS system crashed today, resulting in the loss of many passes (not loss of data, just loss of downlink time). It is hoped that this problem will be fixed today or tomorrow. Extra Weilheim passes are being added in the meantime. The memory fill hit a max of 32% and is at 29% as of the last pass (~20:30 UT), so we are not in danger of losing data, but no actions should be taken that are risky from a memory perspective until the station problems are fixed.

Spacecraft Status

Temperatures are trending normally, with the cold plate temperature 106K and 104.8K (down from last week, normal according to daylight cycle). The CPC temperature is slightly up, to 49.7C. Keep an eye on that one.

Data Gaps

There were several data gaps caused by the Wallops station problems. Most, if not all, of the monitor rate data have been replayed, but here's a list of all the gaps recorded:

 2012-12-17T19:10:00.000 -- 2012-12-17T19:15:00.000       300.00000
 2012-12-16T14:20:00.000 -- 2012-12-16T14:25:00.000       300.00000
 2012-12-16T16:00:00.000 -- 2012-12-16T16:05:00.000       300.00000
 2012-12-16T17:40:00.000 -- 2012-12-16T17:45:00.000       300.00000
 2012-12-16T19:15:00.000 -- 2012-12-16T19:25:00.000       600.00000
 2012-12-16T21:00:00.000 -- 2012-12-16T21:05:00.000       300.00000
 2012-12-15T17:45:00.000 -- 2012-12-15T17:50:00.000       300.00000
 2012-12-15T19:25:00.000 -- 2012-12-15T19:35:00.000       600.00000
 2012-12-15T21:05:00.000 -- 2012-12-15T21:10:00.000       300.00000
 2012-12-15T22:50:00.000 -- 2012-12-15T22:55:00.000       300.00000
 2012-12-14T17:55:00.000 -- 2012-12-14T18:05:00.000       600.00000
 2012-12-14T19:40:00.000 -- 2012-12-14T19:45:00.000       300.00000
 2012-12-14T21:15:00.000 -- 2012-12-14T21:20:00.000       300.00000
 2012-12-14T22:55:00.000 -- 2012-12-14T23:00:00.000       300.00000
 2012-12-13T21:25:00.000 -- 2012-12-13T21:30:00.000       300.00000
 GAP START TIME              GAP END TIME                   GAP (SEC)
 012-12-15T21:05:00.000 -- 2012-12-15T21:10:00.000       300.00000
 GAP START TIME              GAP END TIME                   GAP (SEC)
 2012-12-15T03:45:00.000 -- 2012-12-15T03:50:00.000       300.00000
 2012-12-15T07:20:00.000 -- 2012-12-15T07:45:00.000       1500.0000
 2012-12-14T01:40:00.000 -- 2012-12-14T01:50:00.000       600.00000
 2012-12-14T07:05:00.000 -- 2012-12-14T07:15:00.000       600.00000
 2012-12-14T17:50:00.000 -- 2012-12-14T17:55:00.000       300.00000
 2012-12-14T21:05:00.000 -- 2012-12-14T21:10:00.000       300.00000
 2012-12-12T00:50:00.000 -- 2012-12-12T00:55:00.000       300.00000

Detector issues

Difficulties with detector 8 persisted from last week. (See Tohban Report 2012-12-12.) To summarize, last week D8 events were turned off to ease up the memory, and the voltage was lowered for this detector by 10 steps to attempt to reduce the noise. Last week it was also noticed that D4 was occasionally segmenting, so the HV was *raised* by 5 steps to try and achieve consistent segmentation.

Both efforts were steps in the right direction, though neither problem was completely solved. Noise in D8 was not so long-lasting but was still significant after the voltage stepdown, and so the voltage was lowered by 10 steps on Dec. 14 and again on Dec. 17. The assessment of performance since the Dec. 17 change is being made difficult by the ground station difficulties and temporary apollo server problems at Berkeley; this will be updated a soon as possible. The first data coming back indicate that D8 is still noisy.

D4 segmentation was more consistent after the voltage raise last week, but the detector still unsegmented during the SAA and shortly afterward. The voltage was then raised by 5 steps on Dec. 17. The effect also cannot be fully assessed yet. First data coming back indicate that the segmentation is indeed more stable but is still slipping during and sometimes shortly after the SAA. Fast rates in the rear segment are high whenever the detector is unsegmented; it is thought that this is not a problem.

For the meantime, D8 events remain off, and all events are off at night except for +/-4 minutes.

Here is a summary of all high voltage changes this week (and some from last week):

 (last week) 2012-345-21:39:06 D8 HV lowered by 10 steps
 (last week) 2012-345-21:41:00 D4 HV raised by 5 steps
 2012-349-19:26:12 D8 HV lowered by 10 steps (now 2867.7V)
 2012-352-17:22:36 D8 HV lowered by 10 steps (now 2617.6 V)
 2012-352-17:23:28 D4 HV raised by 5 steps (now 4215.7 V)

The next tohban should monitor D8 and D4 carefully. Further changes to the HV may be necessary. (See Tohban Report 2012-12-12 for more discussion.) If/when the D8 noise appears to be under control, events should be re-enabled for this detector, keeping an eye on the memory fill afterward. If the memory fill is not in danger, then the decimation pattern can be changed back to normal/vigorous. (And events could be taken at night if desired.) Some action on the occasional raised rates in detector 2 may also be necessary since this is cutting into the livetime.

Decimation active/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 min
Night time data (rears) +/- 4 min
Require extra passes? Yes
Requirement for moving pointer? Some data replayed
Attenuator operation Normal
Detector problems? D4 and D8
Personal tools